STRUCTURE OF HIGH-ORDER WATER CLUSTERS OF β-PROPIOLACTONE BY BROADBAND MI-CROWAVE SPECTROSCOPY

JUSTIN L. NEILL, University of Michigan. Department of Astronomy 830 Dennison Bldg. 500 Church St. Ann Arbor, MI 48109-1042; \underline{CRISTOBAL PEREZ}, MATT T. MUCKLE, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904; ISABEL PENA, JUAN C. LOPEZ, JOSE L. ALONSO, Grupo de Espectroscopía Molecular (GEM). Edificio Quifima, Laboratorios de Espectroscopía y Bioespectroscopía. Parque Científico Uva Universidad de Valladolid. 47011 Valladolid, Spain.

The rotational spectra of β-propiolactone-(H_{2}O)_{n} (n=1-5) adducts have been extensively analyzed by broadband microwave spectroscopy (CP-FTMW). Unambiguous identification of their structures has been achieved from the spectra of the parent species and H_{2}^{18}O single substitution clusters and Stark effect measurements. In addition to our previous work, the substitution structures for two n=4 and one n=5 complexes are presented. The three structures show a cyclic arrangement in the oxygen framework. For both n=4 structures, the water molecules form a quasiplanar ring that sits above the BPL unit. The structural differences between these two n=4 complexes are discussed in terms of the water oxygen atom positions and dipole moment orientations, enabling to distinguish between isomers with certainty by CP-FTMW spectroscopy. For n=5, a cyclic arrangement similar to n=4 was found with water molecules making a puckered five-water ring. Substitution coordinates for the oxygen framework support the assignment. Our results show that complexation with BPL induces measurable structural changes in the (H_{2}O)_{n} (n = 3, 4, 5) pure water clusters. This fact is also discussed in terms of the variation in O-O distances within pure and complexed water clusters.